OpenStudy (anonymous):

Don't know where I'm going wrong :S Question: |x-2| = 7 - x My working out: Case 1: x - 2 = 7 - x 2x = 9 x = 9/2 x = 4.5 Case 2: -(x-2) = 7 - x -x + 2 = 7 - x Undef. ∴x = 4.5 Test x = 4.5: |4.5 - 2| = 7 - 4.5 2.5 = 3.5

OpenStudy (angela210793):

how abt.... \[|x-2|=\sqrt{(x-2)^2}\]right?

OpenStudy (anonymous):

Yeah, but how would I make the inequation equal on both sides? :/

OpenStudy (anonymous):

7 - 4.5 = 2.5!!!

OpenStudy (angela210793):

I think only 4.5 is the solution..

OpenStudy (anonymous):

you can draw the graphs ... it will instersect at one and only point i.e. 4.5

OpenStudy (anonymous):

OH GOD, THANK YOU SOTARANTINO, what's the point of doing such questions if I make a mistake at the most basic parts -_- Thanks a lot guys :p,

OpenStudy (angela210793):

hmmm... 7 - 4.5=2.5

OpenStudy (angela210793):

where were my eyes!!!! :@@@@@@@@@@@@

OpenStudy (anonymous):

lol!!!